# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v4

@@ VI Communication optimisation
$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (42a)
${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (42c)
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (42d)
$P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.$ (42e)
Note, that the radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow. To solve this optimisation problem, we must derive the gradients of the objective function and the constraints. Let’s define the objective function as follows,
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\quad\quad\quad\quad\forall j$ (44)
$\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ =$=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (45)
@@
$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vspace{-0.05 in}$ (47)
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vspace{-0.05 in}$ (48)
@@
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vspace{-0.03 in}$ (49)
@@
Solving the optimisation problem defined in equations (42a)-(42d) uses the IPM, as detailed in Algorithm 2. This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This continues until $\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{W}_{\mathrm{c}}^{(n)}\|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{\mathrm{c}}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
@@ VII Integrated Waveform Optimisation for ISAC
$\mathcal{P}_{1}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\quad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (50a)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t\in T,$ (50b)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k\in K,$ (50c)
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (50d)
$P_{\mathrm{T}}=P_{\mathrm{r}}+P_{\mathrm{c}}.$ (50e)
The total power constraint in (50d) limits the combined radar and communication power to $P_{\mathrm{T}}$. To solve this non-convex problem, the ADMM is employed. This approach unifies the projected gradient method used for radar waveform optimisation and the IPM technique used for communication waveform optimisation, efficiently handling the coupling between subsystems while decomposing the complex problem into more manageable subproblems. This involves introducing auxiliary variables $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ such that $\mathbf{Z}_{\mathrm{r}}=\mathbf{W}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}=\mathbf{W}_{\mathrm{c}}$. The augmented Lagrangian for the problem is given as,
@@
$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),$ (51)
@@
$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\quad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (54)
@@
$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).$ (56)
@@
$\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\quad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (57)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t,$ (58)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k,$ (59)
$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},$ (60)
@@
$\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (61)
$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).$ (62)
@@
$\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (63)
@@
$\quad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg{]}.$ (64)
@@
0:  Optimized waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
1:  Initialize $n=0$
@@ VIII-1 Projected Gradient Method with Penalty
$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),$ (65)
@@ VIII-2 Gradient-assisted IPM
$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),$ (66)
@@ VIII-3 ADMM
$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),$ (67)
@@ IX Numerical Results
In this section, we present simulation results for our conventional benchmarks and the three previously introduced optimisation techniques. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section. The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets and $K=3$ UEs are deployed. A pathloss exponent of $\eta=3$ is considered to model large-scale fading. The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters. The channel variance is fixed at $\sigma_{h}^{2}=1$, and the probability of false alarm $P_{FA}$ is set to $10^{-4}$. $\overline{C}_{k}$ is set to $210^{-}$ 2 which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits.
IX-A Trade-off for the conventional benchmark
@@
Figure 1: The ISAC system performance at $P_{\mathrm{r}}=0.5$. a) Radar subsystem CIC design performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem ZF precoding performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
[…]

### v5

@@ VI Communication optimisation
$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\qquad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (45a)
[…]
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (45b)
$P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.\vskip-3.61371pt$ (45c)
Similar to the radar waveform optimisation, fixed power allocation is used between subsystems, where radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow. To solve this optimisation problem, we must derive the gradients of the objective function and the constraints. Let’s define the objective function as follows,
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\qquad\qquad\qquad\qquad\forall j$ (47)
$\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ $=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (48)
@@
$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vskip-3.61371pt$ (50)
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vskip-3.61371pt$ (51)
@@
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vskip-2.168pt$ (52)
@@
Solving the optimisation problem defined in equations (45a)-(45b) uses the IPM, as detailed in Algorithm 2. This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This continues until $\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{W}_{\mathrm{c}}^{(n)}\|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{\mathrm{c}}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
VI-A Communication waveform BER-optimisation
We formulate another baseline design that minimises the instantaneous BER with respect to $\mathbf{W}_{\mathrm{c}}$, while ensuring a minimum radar KLD per target. Where the instantaneous BER of user $k$ is approximated by (54). The optimisation problem is,
$\mathcal{P}_{5}:\min_{\mathbf{W}_{\mathrm{c}}}\qquad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{BER}_{k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (53a)
[…]
$\mathrm{BER}_{k}\;\leq\;\overline{C}_{k},\forall k\in K,$ (53b)
where the radar subsystem precoder $\mathbf{W}_{\mathrm{r}}$ uses the identity covariance $\mathbf{R}_{w}=\mathbf{I}_{N}$. As such, it is important to define the BER, where for $M$-ary constellations, BER of user $k$ is approximated by [52],
$\mathrm{BER}_{k}~\approx~\frac{\lambda_{\mathrm{c}}}{M\,\log_{2}M}\,Q\!\Bigl(\sqrt{\beta_{\mathrm{c}}\,\mathrm{SINR}_{k}}\Bigr),$ (54)
where $\lambda_{\mathrm{c}}$ and $\beta_{\mathrm{c}}$ are constellation-dependent constants (e.g., $\lambda_{\mathrm{c}}=4$, $\beta_{\mathrm{c}}=2$ for QPSK). Unlike the unified closed-form gradients in our KLD-based design, this BER objective depends on the $Q(\cdot)$ function in (54), which introduces additional computational complexity in the optimisation process due to its non-elementary nature and the challenges in evaluating its derivatives efficiently. This distinction illustrates the enhanced analytical tractability of the proposed KLD-based approach. This problem is inherently non-convex and nonlinear due to the $Q$-function in (54) and $\mathrm{KLD}_{\mathrm{r},t}$ constraint. Using $\nabla_{\mathbf{W}_{\mathrm{c}}}\,\mathrm{BER}_{k}$ derived in Appendix C, and the constraint gradients derived in (50)-(52), we employ the same gradient-assisted IPM as in Algorithm 2, to handle the nonconvexities in (53), leveraging the reduced dimensionality of $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$.
@@ VII Integrated Waveform Optimisation for ISAC
$\mathcal{P}_{6}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\qquad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (55a)
[…]
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (55b)
$P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}.$ (55c)
The total power constraint in (55b) limits the combined radar and communication power to $P_{\mathrm{T}}$, and optimises the power allocation between the two subsystems. To solve this non-convex problem, the ADMM is employed. This approach unifies the projected gradient method used for radar waveform optimisation and the IPM technique used for communication waveform optimisation, efficiently handling the coupling between subsystems while decomposing the complex problem into more manageable subproblems. This involves introducing auxiliary variables $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ such that $\mathbf{Z}_{\mathrm{r}}=\mathbf{W}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}=\mathbf{W}_{\mathrm{c}}$. The augmented Lagrangian for the problem is given as,
@@
$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),\vskip-3.61371pt$ (56)
@@
$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\qquad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (59)
@@
$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).\vskip-3.61371pt$ (61)
@@
$\vskip-3.61371pt\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\qquad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (62)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\qquad\forall t,$ (63)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\qquad\forall k,$ (64)
$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},\vskip-3.61371pt$ (65)
@@
$\vskip-3.61371pt\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (66)
$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).\vskip-3.61371pt$ (67)
@@
$\vskip-3.61371pt\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (68)
@@
$\qquad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg].\vskip-3.61371pt$ (69)
@@
0: Optimised waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
1: Initialise $n=0$
@@ VIII-1 Projected Gradient Method with Penalty
$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),\vskip-3.61371pt$ (70)
@@ VIII-2 Gradient-assisted IPM
$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),\vskip-3.61371pt$ (71)
@@ VIII-3 ADMM
$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),\vskip-3.61371pt$ (72)
@@ IX Numerical Results
In this section, we present simulation results for our conventional benchmarks and the three previously introduced optimisation techniques. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section. The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets and $K=3$ UEs are deployed. A pathloss exponent of $\zeta=3$ is considered to model large-scale fading. The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters. The channel variance is fixed at $\sigma_{h}^{2}=1$, the probability of false alarm $P_{FA}$ is set to $10^{-4}$, $\overline{C}_{k}$ is set to $10^{-}2$ which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits, $N_{o}$ is the noise power spectral density, and $P_{\mathrm{R}}$ is the total received power.
IX-A Conventional benchmark
@@
Figure 1: ISAC performance at $P_{\mathrm{r}}=0.5$. a) Radar subsystem CIC design performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem ZF precoding performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
Fig.1 presents the ISAC system performance, analysing both its radar and communication subsystems at $P_{\mathrm{c}}=0.5$. The analysis encompasses the radar subsystem’s target detection capabilities and the communication subsystem’s data transmission quality. Fig.1.a presents the performance of the radar subsystem, illustrated in terms of the $\mathrm{KLD_{r}}$ and the probability of detection $P_{D}$ against $P_{\mathrm{R}}/N_{o}$ for three potential targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$. The curves for all three targets exhibit a consistent upward trend with increasing $P_{\mathrm{R}}/N_{o}$, indicating that higher $P_{\mathrm{R}}/N_{o}$ levels enhance the radar subsystem’s capability to distinguish and detect targets. It can be observed from the figure that at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\mathrm{KLD_{\mathrm{r}}}$ values for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are approximately $\{8.43,\,6.77,\,9.1\}$ bits, respectively, while the corresponding detection probabilities $P_{D}$ reach $\{0.197,\,0.143,\,0.221\}$. By the time $P_{\mathrm{R}}/N_{o}$ increases to $20$ dB, the $\mathrm{KLD_{\mathrm{r}}}$ improves to around $\{19.87,\,17.59,\,20.71\}$ bits, and $P_{D}$ increases to $\{0.771,\,0.627,\,0.803\}$. Throughout these $P_{\mathrm{R}}/N_{o}$ values, $\mathrm{T}_{3}$ remains the most detectable target, consistently demonstrating the highest $\mathrm{KLD_{\mathrm{r}}}$ and $P_{D}$, followed by $\mathrm{T}_{1}$ and then $\mathrm{T}_{2}$. The consistent rise in both $\mathrm{KLD_{\mathrm{r}}}$ and $P_{D}$ with increasing $P_{\mathrm{R}}/N_{o}$ demonstrates the strong correlation between the two measures, confirming the validity of utilising $\mathrm{KLD_{\mathrm{r}}}$ as a reliable measure of radar performance, as well as a reliable design tool. Higher $\mathrm{KLD_{\mathrm{r}}}$ values are directly translated to improved detection probabilities, reinforcing that $\mathrm{KLD_{\mathrm{r}}}$ effectively captures the system’s ability to distinguish targets. Fig.1.b shows the performance of the communication subsystem under the ZF precoding scheme, where the $\mathrm{KLD_{\mathrm{c}}}$ and the BER are plotted versus $P_{\mathrm{R}}/N_{o}$ for three single-antenna UEs. It can be seen that, as $P_{\mathrm{R}}/N_{o}$ increases, the KLD for all UEs rises, reflecting enhanced signal distinguishability. Correspondingly, the BER decreases, indicating better communication quality. Around $10$ dB, however, both an upper bound in KLD and an error floor in BER become evident, highlighting the persistent radar interference as the dominant performance limiting factor once IUI has been largely cancelled by ZF at $P_{\mathrm{c}}=0.5$. Thereby, this validates the use of KLD as an effective tool for characterising and designing communication systems since higher KLD values systematically align with lower BER.
Figure 2: Radar subsystem CIC design performance comparison between communication-assisted and non-assisted configurations at $P_{\mathrm{r}}=\{0.1,0.9\}$: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
Fig. 2 compares communication-assisted and non-assisted radar configurations for $P_{\mathrm{r}}=0.9$ and $P_{\mathrm{r}}=0.1$ for target $\mathrm{T}_{1}$. The assisted configuration exploits communication signals as an additional information source for radar detection, while the non-assisted baseline considers only dedicated radar signals. Importantly, the communication subsystem performance remains unaffected by the choice of assisted or non-assisted radar operation, as the exploitation occurs solely at the radar receiver through processing the reflected signals. For $P_{\mathrm{r}}=0.1$, the communication-assisted configuration demonstrates substantial performance gains. At $P_{\mathrm{R}}/N_{o}=20$ dB, the assisted system achieves $\mathrm{KLD}_{\mathrm{r}}=19.489$ bits compared to $3.842$ bits for the non-assisted case, representing a remarkable improvement. The corresponding detection probabilities reach $P_{D}=0.749$ and $0.086$, respectively. This significant enhancement stems from the high communication power $P_{\mathrm{c}}=0.9$ providing substantial additional signal energy that, when reflected from target $\mathrm{T}_{1}$, contributes to the detection process. At $P_{\mathrm{r}}=0.9$, both configurations show improved performance, with the assisted system maintaining a consistent advantage. At $P_{\mathrm{R}}/N_{o}=20$ dB, the assisted configuration achieves $\mathrm{KLD}_{\mathrm{r}}=19.873$ bits versus $17.795$ bits for the non-assisted scenario, yielding a modest but good improvement. The detection probabilities are $P_{D}=0.771$ and $0.694$, respectively. The reduced relative gain is expected as the dedicated radar signal already provides substantial detection capability, diminishing the proportional contribution from communication signals. The consistent performance enhancement across all SNR levels confirms that our communication-assisted approach effectively exploits the inherent synergy between communication and radar signals.